# Graph Signal Processing for Global Stock Market Volatility Forecasting

### v1

@@ Abstract
The interconnectedness of global financial markets has brought increasing attention to modeling volatility spillover effects. Via incorporating Graph Signal Processing techniques, a novel multivariate framework, extending the traditional Heterogeneous Auto-Regressive model, is developed in the spectral domain constructed by the graph Fourier transformation method. Further, a set of convolution filters with learnable weights is employed to more flexibly aggregate the past mid-term and long-term information. Using 24 global stock market indices, the effectiveness of the proposed model is demonstrated through comprehensive empirical evaluations.
Keywords: realized volatility forecasting; multivariate time series forecasting; spatio-temporal analysis; graph signal processing.
@@ 1 Introduction
In the complex ecosystem of global financial markets, understanding and forecasting volatility is important for investors, risk managers and policymakers. Unlike the retrospective approach of relying on historical volatility, forecasting volatility provides forward-looking estimates that are more reflective of the present market dynamics and potential upcoming changes. Recent studies have emphasized the volatility spillover phenomenon, which refers to the transmission of volatility shocks from one market to others globally (Kanas, 2000; Forbes and Rigobon, 2002; Poon and Granger, 2003; Diebold and Yilmaz, 2009; Yang and Zhou, 2017; Bollerslev et al., 2018). This phenomenon has gained much attention, especially after the highly volatile 2008 global financial crisis (GFC) and the recent COVID-19. The volatility spillover effect highlights the intricacies of the international financial ecosystem, which necessitates a comprehensive framework to consider the relationships between stock markets when conducting volatility forecasting.
This research aims to advance the field of volatility forecasting by developing a multivariate model that integrates the volatility spillover effect across multiple stock markets, via employing the Graph Signal Processing (GSP) technique (Chen et al., 2015; Wu et al., 2021). GSP extends traditional signal processing to data defined on graphs. A graph can be defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes representing the entities in the graph and $\mathcal{E}$ is the set of edges representing the relationships or links between nodes. GSP provides a powerful framework for analyzing and manipulating complex networks and relational data. This study adapts the GSP tools such as the Graph Fourier Transformation (GFT) and spectral analysis techniques to volatility data. The proposed framework and the empirical findings are expected to help financial practitioners better understand the intricate behaviors of global market volatility.
@@
The organization of this paper is as follows. Section 2 provides a review of pertinent literature. Section 3 discusses the motivation for this study. Section 4 details the methods and practical application of the proposed model. Section 5 presents the conducted experiments and their results and the corresponding codes are available at https://github.com/MikeZChi/GSPHAR.git. Section 6 concludes the discussion.
2 Literature Review
This review section aims to provide a clear understanding of how volatility is measured and to examine different univariate and multivariate models used in forecasting volatility.
2.1 Volatility Measurement
Although volatility has been regarded as an unobservable and latent factor behind the fluctuations in an asset’s returns over a given time period, various methods for measuring volatility have been suggested in the literature (Poon and Granger, 2003). Among different measurement approaches, the high-frequency-data-based Realized Volatility (RV), proposed by Andersen and Bollerslev (1998), is used by this research to represent the volatility of different stock markets. The RV assesses the volatility of an asset at time $t$ by tracking its return over a sample period using a sufficiently high frequency of sampling. This measure is particularly useful for capturing the day-to-day fluctuations in a market index and is extensively used in financial practice. Hence, RV is the forecast target of the proposed model and is reviewed in this section.
With access to high-frequency data, RV represents the daily volatility for day $t$ because it is derived from the sum of squared high-frequency intraday returns, as defined below:
$\text{RV}_{t}=\sum_{m=1}^{M}r_{m,t}^{2},$ (1)
where RVt is the RV data on day $t$, the sampling frequency is $\frac{1}{M}$ and $r_{m,t}$ denotes the $m$-th intraday high-frequency return on day $t$.
2.2 Univariate and Multivariate Volatility Forecasting
Since the proposed model is designed to forecast RV time series data, the discussion focuses on models pertinent to RV forecasting problems. Besides, given that the experiments of this research involve forecasting future RV values for stock market indices worldwide, models reviewed in this section target RV forecasting on stock market indices by default. Notably, the Heterogeneous Auto-Regressive (HAR) model, initially introduced by Corsi (2009) and later refined by Bollerslev et al. (2018), stands out as a frequently employed model in RV forecasting. In addition to the univariate HAR model, this section also explores several multivariate extensions of the HAR model.
The univariate HAR model uses the pooled panel data as the pattern of volatility in three previous time intervals to forecast future RV values. The pooled panel data include the past daily RV observation, the average of the past weekly RV values and the average of the past monthly RV values, which represent the short-term, mid-term and long-term volatility behaviors, respectively. Here, RVt denotes the daily RV data of the chosen stock market index on day $t$ calculated as per Equation (1). The univariate HAR model is formulated as:
$\text{RV}_{t}=\alpha+\beta_{d}\text{RV}_{t-1}+\beta_{w}\text{RV}_{t-5:t-1}+\beta_{m}\text{RV}_{t-22:t-1}+\epsilon_{t},\;\;\epsilon_{t}\sim N(0,\sigma^{2}),$ (2)
where $\text{RV}_{t-n:t-1}=\frac{1}{n}\sum_{j=t-n}^{t-1}\text{RV}_{j}$ is the average RV of the stock market index from time $t-1$ to time $t-n$. Here, coefficients $\beta_{d}$, $\beta_{w}$ and $\beta_{m}$ are scalars and they measure how the past daily, weekly and monthly volatility patterns influence future RVs. The intercept term $\alpha$ is also a scalar. Since the original RV data usually shows leptokurtosis (fat tails) and skewness, the log-transformed and square root versions of the RV data are often utilized to improve the forecasting accuracy of the HAR model. However, the univariate HAR model cannot capture the comovement or correlations of the volatility in different stock markets, which significantly limits its ability to deliver more accurate forecasts.
In order to integrate the volatility spillover effect to generate more accurate forecasts, the Vector HAR (VHAR) and the HAR Kitchen Sink (HAR-KS) models are proposed by Bubák et al. (2011) and Liang et al. (2020), respectively. These models align with the general design of the univariate HAR model. They also use past daily RV data as the short-term volatility pattern, past weekly RV data as the mid-term volatility pattern and past monthly RV data as the long-term volatility pattern to forecast future RV values. The VHAR model incorporates lagged daily, weekly and monthly RV data from all other stock market indices to capture the volatility inter-relationships. In contrast, the HAR-KS model only relies on lagged daily RV observations from all indices. The formulation of the VHAR model and the HAR-KS model are presented in Equation (3) and Equation (4), respectively,
$\begin{split}\left(\text{RV}_{i,t}\right)^{1/2}=&\;\beta_{i,0}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,d}\left(\text{RV}_{k,t-1}\right)^{1/2}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,w}\left(\text{RV}_{k,t-5:t-1}\right)^{1/2}\\ &\;+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,m}\left(\text{RV}_{k,t-22:t-1}\right)^{1/2}+\varepsilon_{i,t},\end{split}$ (3)
$\begin{split}\left(\text{RV}_{i,t}\right)^{1/2}=&\;\beta_{i,0}+\sum_{k\in\{1,\ldots,N\}}\beta_{i,k,d}\left(\text{RV}_{k,t-1}\right)^{1/2}+\beta_{i,w}\left(\text{RV}_{i,t-5:t-1}\right)^{1/2}\\ &\;+\beta_{i,m}\left(\text{RV}_{i,t-22:t-1}\right)^{1/2}+\varepsilon_{i,t},\end{split}$ (4)
where both $i$ and $k$ are indices of different stock market indices taken into consideration. Here, $i$ is used to denote the target stock market indices and $k$ indexes the included lagged RV variables from all stock market indices to account for the volatility spillover effect. The square root lagged daily, weekly or monthly RV values from other stock market indices, apart from the target index, are treated as additional attributes to capture the volatility spillover effect.
The VHAR and HAR-KS models not only inherit the simple architecture of the univariate HAR model, but are also empirically proven to outperform it in multivariate RV forecasting problems. However, they have some common limitations. The first is scalability. When the number of relevant stock market indices $N$ increases, the number of parameters of the VHAR and the HAR-KS models grows quadratically. Thus, both models have insufficient scalability. In addition, the volatility spillover effect can be much more complex (e.g., nonlinear) than the linear combinations of past RV data, which cannot be captured by the linear VHAR and HAR-KS models.
Without dramatically sophisticating the HAR model, the Graph Neural Network (GNN) HAR model is proposed by Zhang et al. (2023) to generate more accurate multivariate RV forecasts. Here, the model is denoted as the GNNHAR model. The GNNHAR model precomputes the relational network which remains fixed during the training process. The precomputed relational network can be represented by a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as mentioned in the Introduction section. In graph theory, the structure or connectivity of the relational graph with $N$ nodes can be captured by the adjacency matrix $\mathbf{A}=[\mathbf{A}_{ij}]\in\mathbb{R}^{N\times N}$. For a group of $N$ stock markets, Zhang et al. (2023) used the following way to construct a graph. If stock market $i$ and stock market $j$ are found to be correlated with each other, an edge is constructed to link node $i$ and node $j$, and $\mathbf{A}_{ij}=\mathbf{A}_{ji}=1$. Otherwise, there is no edge between node $i$ and node $j$, and $\mathbf{A}_{ij}=\mathbf{A}_{ji}=0$. Correspondingly, the degree matrix $\mathbf{D}$ can be calculated as: $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{N})$, where $d_{i}=\sum_{j=1}^{N}\mathbf{A}_{ij}$.
GNN models can be divided into two categories: spectral and spatial GNNs. Spectral GNNs rely on the spectral graph theory, whereas spatial GNNs depend on message-passing mechanisms (Wu et al., 2021). In the GNNHAR model, spatial GNNs are applied. Spatial GNNs are specialized neural networks for graph data. It captures the complex interactions between nodes through neighborhood information aggregation and transformation. A neighborhood of a specific node refers to the set of nodes directly connected to the particular node. In each layer, the spatial GNN updates the attributes of a specific entity (node) based on the information of its neighborhood. By stacking multiple layers with linear and nonlinear information transformation functions, each entity in the graph can receive processed information from a larger neighborhood. Suppose $\mathbf{v}_{t}$ contains the square root RV value of each of $N$ stock market indices at day $t$: $\mathbf{v}_{t}=[\left(\text{RV}_{1,t}\right)^{1/2},\left(\text{RV}_{2,t}\right)^{1/2},\ldots,\left(\text{RV}_{N,t}\right)^{1/2}]$ and $\mathbf{v}_{t-n:t-m}$ is the average-pooled data: $\mathbf{v}_{t-n:t-m}=\frac{1}{n-m}\sum_{j=t-n}^{t-m}\mathbf{v}_{t}$. The GNNHAR model adds a $l$-layer spatial GNN component to the HAR model as an additional input variable to capture the volatility spillover effect. The model can be formulated as:
$\begin{split}\mathbf{H}^{(0)}=&\;[\mathbf{v}_{t-1},\mathbf{v}_{t-5:t-2},\mathbf{v}_{t-22:t-6}],\\ \mathbf{H}^{(1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(0)}\mathbf{W}^{(1)}),\\ &\ldots,\\ \mathbf{H}^{(l+1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l+1)}),\\ \mathbf{v}_{t}=&\;\bm{\alpha}+\beta_{d}\mathbf{v}_{t-1}+\beta_{w}\mathbf{v}_{t-5:t-2}+\beta_{m}\mathbf{v}_{t-22:t-6}\\ &\;+\bm{\gamma}\mathbf{H}^{(l+1)}+\bm{\epsilon}_{t}.\end{split}$ (5)
Here, $\mathbf{W}^{(l)}$ is the trainable graph information transformation parameter of layer $l$, representing the linear information transformation, and ReLU()($\cdot$)( ⋅ ) function transforms the graph information nonlinearly. $\bm{\gamma}$ is also a trainable parameter to measure the overall influence of related stock market indices. $\beta_{d}$, $\beta_{w}$ and $\beta_{m}$ are scalars measuring the impact of the lagged panel RV observations similar to the HAR model.
Although all the VHAR, HAR-KS and GNNHAR models follow the same general pattern: adding extra input variables to the HAR model to learn the volatility spillover effect and deliver more accurate multivariate RV forecasts, the GNNHAR model uses spatial GNN to capture the relationships between stock market indices in both linear and nonlinear ways. The inclusion of extra GNN inputs equips the GNNHAR model with additional power to capture the nonlinear volatility spillover effect. In addition, the GNNHAR model applies the same set of weights $\mathbf{W}^{(1)}$, $\mathbf{W}^{(2)}$, …, $\mathbf{W}^{(l+1)}$} to different parts of the volatility network to model the volatility spillover effect. This design ensures the GNNHAR model is scalable. Although this seems to bring the GNNHAR model more power to scale up to larger graphs, trying different numbers of layers $l$ to determine the best scope of the volatility spillover influence can be very time-consuming. In addition, the pattern of the volatility spillover effect can vary across time, which reflects different periods of strengthened or weakened correlations (Diebold and Yilmaz, 2012; Antonakakis et al., 2018; Gong et al., 2021). Consequently, the impact range of the volatility spillover effect can also change over time. Thus, the fixed number of layers ($l$) in the GNNHAR model can limit its adaptability and generalizability.
3 Research Motivation
Even though the univariate HAR model and its multivariate extensions have their own constraints, the general architecture of the HAR model is still preferred in the literature. The HAR model is designed to account for the fact that different market participants (such as high-frequency traders, institutional investors, and long-term investors) operate on different time horizons. The HAR model incorporates pooled RV data from these various time scales, typically daily, weekly and monthly. This model design helps to capture the heterogeneous nature of volatility patterns in financial markets. In addition, empirically it achieves good out-of-sample RV forecasting performance with its simple and intuitive structure, which makes it even more attractive to researchers and practitioners.
Based on the discussions in the previous section, the main challenges for multivariate HAR models are their insufficient expressiveness and limited scalability to larger financial systems. These limitations stem from their architectures where each multivariate HAR model tries to account for the volatility spillover effect by including extra input variables in the model. Although the GNNHAR model attempts to address these issues by applying spatial GNNs, the fixed volatility spillover range reduces effectiveness when modeling complex RV dynamics. In addition, incorporating extra input variables to model volatility spillover effectively learns the residuals between the forecast values generated by the univariate HAR model and the true values. However, treating volatility spillover as merely a compensatory mechanism for the univariate HAR model potentially underestimates its significance. According to Poon and Granger (2003), both the interdependent and the autoregressive behaviors of multi-asset volatility dynamics are salient phenomena. Thus, the volatility spillover effect should be as important as the autoregressive relationship inherent in the HAR models, rather than as an additive correction for neglecting cross-entity relationships. The existing multivariate HAR models do not prioritize these volatility inter-relationships, which could potentially limit their overall forecasting capability. Furthermore, the volatility spillover effect is unlikely to remain constant over time, because market conditions, economic factors and investor behavior continuously change. However, few multivariate RV forecasting models adequately account for this inherent dynamic nature.
This research aims to overcome the above constraints and further improve the multivariate RV forecasting accuracy by proposing another multivariate extension of the HAR model. This model relies on GSP techniques, so the model is referred to as the GSPHAR model. The general idea is to build the HAR model in the spectral domain constructed by the GFT method so that the relational information can be well merged with the RV data. Further, a set of convolution filters with learnable weights is employed to more flexibly aggregate the past mid-term and long-term information. The proposed model can also incorporate extra designs to model the dynamic volatility spillover effect. A Demonstration of the proposed framework is shown in Figure 2 in the following methodology section.
4 Methodology and Implementation
Since the proposed model relies on GSP techniques, this section introduces key concepts from GSP and spectral graph theory, with a focus on the GFT and the magnet framework (Zhang et al., 2021; Lin and Gao, 2023). In addition, this section discusses the graph construction method used to model the RV relationship network among different market indices and a simple yet effective dynamic modeling method.
4.1 Spectral Graph Theory and GSP
As briefly discussed in the Introduction and Literature Review sections, a graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the node set and $\mathcal{E}$ is the edge set of the graph. So far we only consider the undirected and unweighted graphs. However, edges between nodes can have attributes such as direction and weights. A directed edge $e_{m}\in\mathcal{E}$ can be formulated as $e_{m}=(v_{\text{start}},v_{\text{end}})$. It starts from node $v_{\text{start}}\in\mathcal{V}$ and ends at node $v_{\text{end}}\in\mathcal{V}$, which points the information flow direction. In contrast, for an undirected edge, information can travel in both directions without any designated starting or end nodes. An undirected edge can also be regarded as a bidirectional edge because $e_{m}=(v_{i},v_{j})=(v_{j},v_{i})$. According to whether the edges are directed or not, graphs can be divided into undirected graphs and directed graphs, as shown in Figure 1 below. Weights are another edge attribute. Weights are usually positive real numbers and they can be interpreted as the importance of the corresponding edges. Depending on whether the edges have weights or not, graphs can be divided into unweighted graphs and weighted graphs.
(a) An example of undirected graph
(b) An example of directed graph
Figure 1: Undirected and directed graphs
@@ 4.1 Spectral Graph Theory and GSP
The adjacency matrix $\mathbf{A}$ is used to capture the structure of graphs. If graph $\mathcal{G}$ has $N$ nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$. For an undirected graph, its adjacency matrix is always symmetric because if $(v_{i},v_{j})\in\mathcal{E}$, $(v_{j},v_{i})\in\mathcal{E}$ as well. Elements in $\mathbf{A}$ can reflect the edge weights. The elements in the adjacency matrix of an undirected and unweighted graph can only take values of 00 or $1$ to reflect connectivity, as shown below:
$\mathbf{A}_{ij}=\begin{cases}1,&\mbox{$(v_{i},v_{j})\in\mathcal{E}$};\\ 0,&\mbox{$(v_{i},v_{j})\notin\mathcal{E}$}.\end{cases}$ (6)
However, for a directed and weighted graph, if $e_{m}=(v_{i},v_{j})\in\mathcal{E}$ and the weights attached to edge $e_{m}$ is $w_{ij}$, $\mathbf{A}_{ij}=w_{ij}$, and in general $\mathbf{A}_{ij}\neq\mathbf{A}_{ji}$. The formulation of the adjacency matrix of the directed and weighted graph is presented below:
@@
No matter whether the graph is directed or weighted, the degree matrix $D$ is always formulated as $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{N})$, where $d_{i}=\sum_{j=1}^{N}\mathbf{A}_{ij}$, for a graph $\mathcal{G}$ with $N$ nodes.
The graph Laplacian matrix $\mathbf{L}$ can be formulated based on the adjacency matrix $\mathbf{A}$ and the degree matrix $\mathbf{D}$:
$\mathbf{L}=\mathbf{D}-\mathbf{A}.$ (8)
The normalized graph Laplacian matrix $\widetilde{\mathbf{L}}$ can be formulated as:
$\widetilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (9)
The degree matrix $\mathbf{D}$ is a diagonal matrix and thus always symmetric, while the adjacency matrix $\mathbf{A}$ is only symmetric when the corresponding graph $\mathcal{G}$ is undirected. If both $\mathbf{D}$ and $\mathbf{A}$ are symmetric and elements in $\mathbf{A}$ (i.e., edge weights) are non-negative, the normalized graph Laplacian matrix $\widetilde{\mathbf{L}}$ is also symmetric, positive semidefinite and diagonalizable. Its eigendecomposition can be formulated as:
$\widetilde{\mathbf{L}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ (10)
where the columns of $\mathbf{U}$ are the eigenvectors of $\widetilde{\mathbf{L}}$, the matrix $\bm{\Lambda}$ is a diagonal matrix containing the corresponding non-negative eigenvalues of $\widetilde{\mathbf{L}}$ and $\mathbf{U}^{\top}$ is the transpose of $\mathbf{U}$. Especially, the eigenvectors in $\mathbf{U}$, which are referred to as the graph Fourier bases, can form a set of orthonormal bases and the eigenvalues are arranged in ascending order from the top left element to the bottom right element in $\bm{\Lambda}$. Large eigenvalues indicate that the corresponding eigenvectors contain high-frequency graph signals, whereas low eigenvalues indicate that the corresponding eigenvectors contain low-frequency graph signals. For simplicity, suppose an undirected graph has $N$ nodes and each node only has a scalar feature so that the graph signal can be represented as a vector $\mathbf{x}\in\mathbb{R}^{N}$. Based on the eigendecomposition of the normalized graph Laplacian matrix of an undirected graph, the GFT and inverse GFT (IGFT) can be formulated in Equations (11) and (12), respectively,
$\widetilde{\mathbf{x}}=\mathbf{U}^{\prime}\mathbf{x};$ (11)
$\mathbf{x}=\mathbf{U}\widetilde{\mathbf{x}},$ (12)
where $\widetilde{\mathbf{x}}$ is the graph signal projected on the spectral domain.
Compared to spatial methods, the spectral graph signal $\widetilde{\mathbf{x}}$ contains more global information. Through GFT, the graph signal $\mathbf{x}$ is aggregated to produce its projection $\widetilde{\mathbf{x}}$ in the spectral domain according to the overall graph structure instead of the neighborhood information only. However, it is challenging to perform such GFT and IGFT based on eigendecomposition on directed graphs, whereas the spatial methods can naturally extend to directed graphs based on their neighborhood information passing design. Fortunately, a unified GSP framework can be applied to both undirected and directed graphs based on the magnetic Laplacian proposed by Zhang et al. (2021).
4.2 Magnetic Laplacian
Although the spatial methods can be more easily adapted to directed graphs compared to the spectral methods, they also have some drawbacks. Relying on the neighborhood information passing mechanism, spatial methods place more emphasis on out-going edges rather than in-going edges, thus limiting their ability to process the information on directed graphs (Zhang et al., 2021). The magnet Laplacian can give equal importance to different directions with its special design.
Suppose a directed graph $\mathcal{G}$ has $N$ nodes. In order to compute the magnet Laplacian $\mathbf{L}_{m}$, the symmetric variant of adjacency matrix $\mathbf{A}^{s}$ and the corresponding degree matrix $\mathbf{D}^{s}$ should be firstly defined in Equation (13) and Equation (14), respectively,

### v2

# Graph Signal Processing for Global Stock Market Realized Volatility Forecasting

@@ Abstract
This paper introduces an innovative realized volatility (RV) forecasting framework that extends the conventional Heterogeneous Auto-Regressive (HAR) model via integrating the Graph Signal Processing (GSP) technique. The volatility spillover effect is embedded and modeled in the proposed framework, which employs the graph Fourier transformation method to effectively analyze the global stock market dynamics in the spectral domain. In addition, convolution filters with learnable weights are applied to capture the historical mid-term and long-term volatility patterns. The empirical study is conducted with RV data of $24$ global stock market indices with around $3500$ common trading days from May 2002 to June 2022. The proposed model’s short-term, middle-term and long-term RV forecasting performance is compared with various HAR type models and the graph neural network based HAR model. The results show that the proposed model consistently outperforms all other models considered in the study, demonstrating the effectiveness of integrating the GSP technique into the HAR model for RV forecasting.
Keywords: volatility spillover; realized volatility; graph signal processing; spectral analysis.
@@ 1 Introduction
In the complex ecosystem of global financial markets, understanding and forecasting volatility is important for investors, risk managers and policymakers. Recent studies on financial volatility analysis have emphasized the volatility spillover phenomenon, which refers to the transmission of volatility shocks from one market to others globally (Kanas, 2000; Forbes and Rigobon, 2002; Poon and Granger, 2003; Diebold and Yilmaz, 2009; Yang and Zhou, 2017; Bollerslev et al., 2018). This phenomenon has gained much attention, especially after the highly volatile 2008 global financial crisis (GFC) and the recent COVID-19. The significance of the volatility spillover effect necessitates a comprehensive volatility forecasting framework to consider both the cross-sectional volatility inter-relationship among different financial assets and the temporal autocorrelation of the volatility of each asset.
Various financial volatility measurements have been suggested in the literature (Poon and Granger, 2003). Among different measurement approaches, the high frequency data based Realized Variance (RV), proposed by Andersen and Bollerslev (1998), is particularly useful for capturing the intraday fluctuations in financial time series and is extensively used in the financial practice. The daily RV is calculated as the sum of squared intraday high frequency returns. The Heterogeneous Auto-Regressive (HAR) model, initially introduced by Corsi (2009) and later extended by numerous studies, such as Patton and Sheppard (2015); Bollerslev et al. (2018), stands out as a frequently employed model in RV forecasting due to its effectiveness and simplicity. The conventional HAR type models cannot capture the volatility spillover effect among different assets. To address this, several HAR model extensions, including the Vector HAR (VHAR) model (Bubák et al., 2011), the HAR Kitchen Sink (HAR-KS) model (Liang et al., 2020) and the Graph Neural Network HAR model (GNNHAR) model (Zhang et al., 2025) are proposed. Meanwhile, all these extensions incorporate the volatility spillover effects into the HAR framework by treating them as exogenous variables, via simultaneously considering the spatial and temporal relationships of volatility dynamics.
With the development of graph signal processing (GSP) technique, which extends the traditional signal processing methods to graphical data, the volatility spillover effect and the volatility auto-regressive process can be analyzed not only in the spatial and temporal domains, but also in the graph spectral domain. Comparing to the existing volatility forecasting studies which are limited to spatial and temporal domain, incorporating the additional graph spectral domain to model the volatility spillover effect can help fully explore the inherent structure of the volatility dynamics and identify both the local and global pattern of the multi-asset volatility interrelationships. However, the current literature lacks explorations of GSP-related strategies to effectively integrate the volatility spillover effect within the HAR framework. Motivated by the identified gap, this research proposes a novel framework based on the GSP technique to embed the volatility spillover effect in the HAR model. The proposed framework is referred to as the GSPHAR. The main contributions of this research are summarized below:
  • The proposed GSPHAR framework is the first attempt to explore a GSP-based strategy to incorporate the volatility spillover effect into the HAR model. The GSPHAR framework does not treat the volatility interrelationships as exogenous variables in the HAR model. Rather, the volatility spillover effect is prioritized and embedded in the design of the proposed framework. Empirical results show that the proposed approach can embed and effectively capture the volatility spillover effects in the HAR framework and significantly enhance the RV forecasting accuracy.
  • Via employing the convolution filters with learnable weights, the proposed GSPHAR framework incorporates more flexible volatility auto-regressive process, comparing to the existing HAR type models. The decaying pattern of the influence from historical volatility to future volatility does not only exist between the short-term, mid-term and long-term periods as the HAR does, but it also occurs at higher granularity within the mid-term and long-term periods.
  • The proposed GSPHAR framework allows extra designs to integrate the changing volatility interrelationships into RV forecasting tasks. These designs allow the proposed framework to capture the dynamic nature of the volatility spillover effect, which is overlooked in the other HAR model extensions mentioned above.
@@
The paper is organized as follows. Section 2 provides preliminaries of pertinent concepts and models. The proposed framework is developed in Section 3. Section 4 presents the conducted experiments and their results. The implementation code is available at https://github.com/MikeZChi/GSPHAR.git. Section 5 concludes the paper.
2 Background
This section reviews relevant concepts on graphs and GSP techniques. In addition, different baseline HAR models used in RV forecasting are reviewed.
2.1 Graph Theory Fundamentals
A graph can be denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the node set and $\mathcal{E}$ is the edge set of the graph. An edge $e_{ij}\in\mathcal{E}$ can be formulated as $e_{ij}=(v_{i},v_{j})$, which indicates that the edge $e_{ij}$ connects node $v_{i}$ and node $v_{j}$. Edges can have attributes such as directions and weights. The direction of an edge defines how information flows between nodes, and the weight of an edge represents the importance or strength of the connection.
@@ 2.1 Graph Theory Fundamentals
The adjacency matrix $\mathbf{A}$ is used to capture the structure of graphs. If graph $\mathcal{G}$ has $N$ nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$. For an undirected graph, its adjacency matrix is always symmetric because if $(v_{i},v_{j})\in\mathcal{E}$, $(v_{j},v_{i})\in\mathcal{E}$ as well. Otherwise, the adjacency matrix of a directed graph is asymmetric. Elements in $\mathbf{A}$ reflect the edge weights. Suppose the weight attached to edge $e_{ij}$ is a positive real number $w_{ij}$. The formulation of the adjacency matrix is presented below:
@@
The degree matrix $D$ is formulated as $\mathbf{D}=\text{diag}(d_{1},\ldots,d_{N})$, where $d_{i}=\sum_{j=1}^{N}\mathbf{A}_{ij}$.
2.2 GSP Based on the Magnetic Laplacian
A unified GSP framework can be applied to both undirected and directed graphs based on the magnetic Laplacian proposed by Zhang et al. (2021).
Suppose a directed graph $\mathcal{G}$ has $N$ nodes. In order to compute the magnetic Laplacian $\mathbf{L}_{m}$, the symmetric variant of adjacency matrix $\mathbf{A}^{s}$ and the corresponding degree matrix $\mathbf{D}^{s}$ should be firstly defined in Equation (2) and Equation (3), respectively,